OpenStudy (anonymous):

Working together, paul and daniel can pick forty bushels of apples in 4.95 hours.had he done it alone it would have taken daniel 9 hours. Find how long it would take paul to do it alone?

hero (hero):

The working together formula for Paul, P, and Daniel, D is: \[\frac{PD}{P + D} = t\] Since t = 4.95 and d = 9, then: \[\frac{P(9)}{P + 9} = 4.95\] Solve for P: 9P = 4.95(P + 9) 9P = 4.95P + 44.55 9P - 4.95P = 44.55 (9 - 4.95)P = 44.55 4.05P = 44.55 P = 44.55/4.05 P = 11

hero (hero):

Thus, working alone, Paul can do it in 11 hour

hero (hero):

@LeaGamage